# myequation

### v1

@@ 3.2 Impact of Bell State Measurement Imperfections
Figure 2 presents the hashing rate achieved by each protocol in the simulation as increasing values of $\lambda_{\text{BSM}}$. Figures 2(d)– 2(f) show the impact of increasing $\lambda_{\text{BSM}}$ on hashing rate for the sequential protocol, while Figures 2(a)– 2(c) display the corresponding results for the parallel protocol.
@@
Figure 2: Comparison of imperfect BSM tolerance in sequential and parallel protocols. Subfigures (a–c) give the hashing rate for the sequential protocol with increasing $\lambda_{\text{BSM}}$. Subfigures (d–f) show results for the parallel protocol. For sufficiently high $\lambda_{\text{BSM}}$, the parallel protocol far more robust to imperfect BSM than the sequential.
@@
In this section, we presented algorithms for simulating the asynchronous distribution of Bell pairs using both sequential and parallel protocols in noisy quantum networks, based on the noise model derived in Section 2. We evaluated the performance of each protocol using our simulations, for networks up to 252525 nodes and quantum memory dephasing times ranging from $10^{-4}$ to $10^{-1}$ seconds. In each experiment we compared the output fidelity of the final distributed Bell pairs, as well as the hashing rate achieved by each protocol. We found that the sequential protocol, although requiring fewer qubits be stored in memory and thereby reducing individual quantum memory decoherence, necessitates a longer overall execution time, which results in alternative decoherence from the extended runtime. In contrast, the parallel protocol utilizes numerous quantum memories simultaneously, but benefits from a significantly faster execution.
@@ 4 Discussion
In this work, we present a lightweight simulation framework for analyzing contemporary asynchronous entanglement distribution protocols in linear quantum networks. We begin by identifying and modeling the dominant sources of error in near-term quantum hardware, which we use to derive a generic and scalable error model for simulating Bell state distribution. One important result from our error analysis is the finding that the fidelity of a Bell state distributed across a quantum network depends only on the total quantum memory time experienced by all qubits, rather than how memory times are distributed among individual qubits. This insight enabled the design of our simplified and efficient simulation algorithms for both sequential and parallel entanglement distribution. We further utilize the results of our error analysis to determine meaningful success metrics for benchmarking the performance of each protocol under realistic noise conditions. Our approach largely improves the practicality and utility of entanglement distribution simulations, offering a valuable tool for assessing protocol viability on noisy intermediate-scale quantum (NISQ) devices.
@@
Using our simulation framework, we evaluate the performance of leading asynchronous entanglement distribution protocols: sequential distribution, where nodes establish entanglement one at a time along the repeater line, and parallel distribution, where all nodes attempt simultaneous and independent entanglement generation. We analyze the performance of each protocol by evaluating the fidelity of the distributed Bell pair, as well as the hashing rate (the rate at which purified Bell pairs can be produced). We simulate and compare each protocol across a range of network sizes and different quantum memory dephasing times, and determine that parallel distribution outperforms sequential distribution in all cases. While final Bell state fidelity was comparable in both protocols, the significantly shorter runtime of the parallel protocol resulted in much higher hashing rates. Moreover, when incorporating the effects of imperfect Bell state measurement we find that both protocols exhibit a strong sensitivity to reduced measurement fidelity. Our findings suggest that parallel asynchronous protocols are particularly well-suited for application in near-term quantum networks.
@@
In this work, we focused on key sources of error in asynchronous entanglement distribution protocols, including optical loss, quantum memory decoherence, and imperfect Bell state measurement. Future extensions of this simulation framework could incorporate additional sources of error, such as detector dark counts and imperfect Bell state preparation. We anticipate these effects can be integrated into our existing noise model through multiplicative corrections to the final fidelity expression in Eq. (17), thereby broadening the scope of errors that can be effectively captured through simple memory time calculations.
@@
Our simulation framework can be directly adapted to study traffic management and routing challenges in multiparty quantum networks. One primary requirement for distributed quantum computation is the ability to distribute GHZ-type entanglement across spatially separated nodes [34, 35, 36, 37, 38]. However, many existing entanglement routing protocols are developed in the absence of realistic noise or hardware constraints. As demonstrated in Appendix A.4, our framework supports the distribution of $n$-qubit GHZ states, enabling an analysis of GHZ-type entanglement routing to be conducted under practical noise conditions. By simulating concurrent entanglement distribution requests across various network topologies, our framework can be used to design and benchmark scheduling algorithms, assess resource consumption, and identify performance bottlenecks. Each of these elements are essential for the successful implementation and scalability of quantum network architectures.
In this work we specifically focus on the noise affecting entanglement distribution in near-term quantum networks. However, our framework can naturally be extended to study the distribution of additional quantum resources across distant nodes a network. One particularly important resource for establishing quantum advantage is quantum magic [39, 40, 41], which captures the degree of non-stabilizerness in a quantum system and correlates directly with the intractability of classically simulating that system. Moreover, recent work has demonstrated that magic can exist non-locally [42, 43], and can be encoded in the entanglement structure of the system. While magic is essential to achieve computational speedups for quantum algorithms, its role in quantum networks remains largely unexplored. In pursuit of understanding non-local magic distribution, we seek to utilize our simulation framework to explore the feasibility and complexity of distributing magic across a quantum network.
The quantum networks studied throughout this paper can be modeled using graph states [34], where vertices represent qubits (or more generally collections of interacting qubits) and edges correspond to entanglement. Numerous techniques have been developed to analyze the generation and dynamics of entanglement in graph states [44, 45, 46, 47, 48], leveraging the underlying algebraic structure to derive exact constraints on entanglement possibilities [49, 50, 51, 52]. In quantum networks, similar constraints naturally emerge from fundamental and practical limitations in entanglement generation and distribution. By combining our realistic simulation model with an algebraic analysis of graph state entanglement we can more accurately probe the potential for noisy quantum networks to achieve desired entanglement arrangements and generation rates [47, 53], offering diagnostics for quantum network capabilities.
@@ A.1 Dephasing Error on a Four Qubit Network
We begin by modeling the effect of dephasing on a two-qubit Bell state, $({|{\phi^{+}}\rangle}\bra{\phi^{+}})_{AB}$, where qubits $A$ and $B$ remain in memory for times $t_{A}$ and $t_{B}$, respectively. The evolution is captured using the following Kraus map:
@@
$\rho_{AB}$ $=\sum^{1}_{i,\,j=0}\;K_{ij}\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}\;K_{ij}^{\dagger}$ (26)
$|\displaystyle=(1-p_{A})\,(1-p_{B})\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}+(1-p_{A})\;p_{B}\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ | (27)
$|\displaystyle+p_{A}\,(1-p_{B})\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}+p_{A}\,p_{B}\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}$ | (28)
@@
$|\rho_{AB}=\alpha_{AB}\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AB}\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ | (30)
@@
$=\big{(}\alpha_{AB}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AB}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}\big{)}\,\otimes\,\big{(}\alpha_{CD}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{CD}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}\big{)}$ (32)
@@
$|\displaystyle=(\alpha_{AB}\,\alpha_{CD}+\beta_{AB}\,\beta_{CD})\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}\,+\,(\alpha_{AB}\,\beta_{CD}+\beta_{AB}\,\alpha_{CD})\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ | (35)
@@ A.2 Generalization to N-Qubit Networks
$=\big{(}\alpha_{AD}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AD}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}\big{)}\,\otimes\,\big{(}\alpha_{EF}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{EF}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}\big{)}$ (38)
@@
[…]
$|\rho_{ADEF}=\big{(}\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}\big{)}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}\,+\,\big{(}\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}\big{)}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ | (39)
@@ A.3 Imperfect Bell State Measurement
$\rho_{ABCD}=(\alpha_{AB}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AB}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}})\,\otimes\,(\alpha_{CD}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{CD}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}})$ (41)
@@
$\rho$ $=\lambda_{BSM}\,\Big{[}(\alpha_{AB}\,\alpha_{CD}+\beta_{AB}\,\beta_{CD}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AB}\,\beta_{CD}+\beta_{AB}\,\alpha_{CD}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big{]}$ (42)
$|\displaystyle\quad+(1-\lambda_{BSM})\,\Big{[}\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$ |
$\hskip 85.35826pt+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big{]}$
@@
$\rho_{1}\otimes\rho_{EF}=(\alpha_{AD}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AD}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}})\otimes(\alpha_{EF}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{EF}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}})$ (44)
@@
$\rho$ $=\lambda_{BSM}^{\prime}\,\Big{[}(\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big{]}$ (45)
$|\displaystyle\quad+\,(1-\lambda_{BSM}^{\prime})\,\Big{[}\alpha_{AD}\,\alpha_{EF}{|{++}\rangle}\bra{++}+\alpha_{AD}\,\beta_{EF}{|{+-}\rangle}\bra{+-}$ |
$\hskip 85.35826pt+\beta_{AD}\,\alpha_{EF}{|{-+}\rangle}\bra{-+}+\beta_{AD}\,\beta_{EF}{|{--}\rangle}\bra{--}\Big{]}$
@@
$\rho_{2}\otimes\rho_{EF}$ $|\displaystyle=\,\Big{[}\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$ | (46)
$+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big{]}$
$\otimes(\alpha_{EF}\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{EF}\,{|{\phi^{-}}\rangle}\bra{\phi^{-}})$
@@
$\rho$ $=\lambda_{BSM}\,\lambda_{BSM}^{\prime}\,\Big{[}(\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big{]}$ (47)
$|\displaystyle\quad+\lambda_{BSM}\,(1-\lambda_{BSM}^{\prime})\,\Big{[}\alpha_{AD}\,\alpha_{EF}{|{++}\rangle}\bra{++}+\alpha_{AD}\,\beta_{EF}{|{+-}\rangle}\bra{+-}$ |
$\hskip 85.35826pt+\beta_{AD}\,\alpha_{EF}{|{-+}\rangle}\bra{-+}+\beta_{AD}\,\beta_{EF}{|{--}\rangle}\bra{--}\Big{]}$
$|\displaystyle\quad+(1-\lambda_{BSM})\,\Big{[}\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$ |
$\hskip 85.35826pt+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big{]}$

### v2

@@ 3.2 Impact of Bell State Measurement Imperfections
Figure 2 presents the hashing rate achieved by each protocol in the analysis as increasing values of $\lambda_{\text{BSM}}$. Figures 2(d)2(f) show the impact of increasing $\lambda_{\text{BSM}}$ on hashing rate for the sequential protocol, while Figures 2(a)2(c) display the corresponding results for the parallel protocol.
@@
Figure 2: Comparisons of imperfect BSM tolerance in sequential and parallel protocols. Protocols are compared in networks with attenuation length $L_{att}=22$km and internode distance L=50/NL=50/Nkm, with varying dephasing times $T_{dp}$ and number of nodes NN. Subfigures (a–c) give the hashing rate for the sequential protocol with increasing $\lambda_{\text{BSM}}$. Subfigures (d–f) show results for the parallel protocol. For sufficiently high $\lambda_{\text{BSM}}$, the parallel protocol is far more robust to imperfect BSM than the sequential.
@@
In this section, we presented algorithms for simulating the asynchronous distribution of Bell pairs using both sequential and parallel protocols in noisy quantum networks, based on the noise model derived in Section 2. We evaluated the performance of each protocol, for networks up to $25$ nodes and quantum memory dephasing times ranging from $10^{-4}$ to $10^{-1}$ seconds. In each experiment we compared the output fidelity of the final distributed Bell pairs, as well as the hashing rate achieved by each protocol. We found that the sequential protocol, although requiring fewer qubits be stored in memory and thereby reducing individual quantum memory decoherence, necessitates a longer overall execution time, which results in alternative decoherence from the extended runtime. In contrast, the parallel protocol utilizes numerous quantum memories simultaneously, but benefits from a significantly faster execution.
@@ 4 Discussion
In this work, we present a compact analytical framework for evaluating contemporary asynchronous entanglement distribution protocols in linear quantum networks. By reducing the simulation of complex quantum processes to simple memory-time calculations, our approach captures the essential physics of noise and decoherence while remaining computationally efficient. We begin by identifying and modeling the dominant sources of error in near-term quantum hardware, which we use to derive a generic and scalable error description of Bell state distribution fidelity. One important result from our error analysis is the finding that the fidelity of a Bell state distributed across a quantum network depends only on the total quantum memory time experienced by all qubits, rather than how memory times are distributed among individual qubits. This insight underpins the simplicity and scalability of our framework for both sequential and parallel protocols, providing an accessible and transparent tool for analyzing entanglement distribution protocols efficiently under realistic noise conditions.
@@
Using our analytical framework, we evaluate the performance of leading asynchronous entanglement distribution protocols: sequential distribution, where nodes establish entanglement one at a time along the repeater line, and parallel distribution, where all nodes attempt simultaneous and independent entanglement generation. We analyze the performance of each protocol by evaluating the fidelity of the distributed Bell pair, as well as the hashing rate (the rate at which purified Bell pairs can be produced). We analyze and compare each protocol across a range of network sizes and different quantum memory dephasing times, and determine that parallel distribution outperforms sequential distribution in all cases. While final Bell state fidelity was comparable in both protocols, the significantly shorter runtime of the parallel protocol resulted in much higher hashing rates. Moreover, when incorporating the effects of imperfect Bell state measurement we find that both protocols exhibit a strong sensitivity to reduced measurement fidelity. Our findings suggest that parallel asynchronous protocols are particularly well-suited for application in near-term quantum networks.
In this work, we focused on key sources of error in asynchronous entanglement distribution protocols, including optical loss, quantum memory decoherence, and imperfect Bell state measurement. In contrast to previous studies that compute fidelity via detailed memory timing histories [27], our framework offers a lightweight analytical approach based solely on total memory time and the number of swapping operations. Expressing the output fidelity solely as a function of total memory time, we reveal the fundamental scaling behavior governing asynchronous protocols. This streamlined formulation enables efficient network analysis at significantly larger qubit scales.
@@
Future extensions of this analytical framework could incorporate additional sources of error, such as detector dark counts and imperfect Bell state preparation. We anticipate these effects can be integrated into our existing noise model through multiplicative corrections to the final fidelity expression in Eq. (17), thereby broadening the scope of errors that can be effectively captured through simple memory time calculations.
@@
Our analytical framework can be directly adapted to study traffic management and routing challenges in multiparty quantum networks. One primary requirement for distributed quantum computation is the ability to distribute GHZ-type entanglement across spatially separated nodes [34, 35, 36, 37, 38]. However, many existing entanglement routing protocols are developed in the absence of realistic noise or hardware constraints. As demonstrated in Appendix A.4, our framework supports the distribution of $n$-qubit GHZ states, enabling an analysis of GHZ-type entanglement routing to be conducted under practical noise conditions. By analyzing concurrent entanglement distribution requests across various network topologies, our framework can be used to design and benchmark scheduling algorithms, assess resource consumption, and identify performance bottlenecks. Each of these elements are essential for the successful implementation and scalability of quantum network architectures.
In this work we specifically focus on the noise affecting entanglement distribution in near-term quantum networks. However, our framework can naturally be extended to study the distribution of additional quantum resources across distant nodes a network. One particularly important resource for establishing quantum advantage is quantum magic [39, 40, 41], which captures the degree of non-stabilizerness in a quantum system and correlates directly with the intractability of classically simulating that system. Moreover, recent work has demonstrated that magic can exist non-locally [42, 43], and can be encoded in the entanglement structure of the system. While magic is essential to achieve computational speedups for quantum algorithms, its role in quantum networks remains largely unexplored. In pursuit of understanding non-local magic distribution, we seek to utilize our analytical framework to explore the feasibility and complexity of distributing magic across a quantum network.
The quantum networks studied throughout this paper can be modeled using graph states [34], where vertices represent qubits (or more generally collections of interacting qubits) and edges correspond to entanglement. Numerous techniques have been developed to analyze the generation and dynamics of entanglement in graph states [44, 45, 46, 47, 48], leveraging the underlying algebraic structure to derive exact constraints on entanglement possibilities [49, 50, 51, 52]. In quantum networks, similar constraints naturally emerge from fundamental and practical limitations in entanglement generation and distribution. By combining our realistic analytical model with an algebraic analysis of graph state entanglement we can more accurately probe the potential for noisy quantum networks to achieve desired entanglement arrangements and generation rates [47, 53], offering diagnostics for quantum network capabilities.
@@ A.1 Dephasing Error on a Four Qubit Network
[…]
@@
[…]
$|\displaystyle=(1-p_{A})\,(1-p_{B})\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}+(1-p_{A})\;p_{B}\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ (28)
$|\displaystyle+p_{A}\,(1-p_{B})\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}+p_{A}\,p_{B}\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}$ (29)
@@
$|\rho_{AB}=\alpha_{AB}\;{|{\phi^{+}}\rangle}\bra{\phi^{+}}+\beta_{AB}\;{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ (31)
@@
[…]
@@
$|\displaystyle=(\alpha_{AB}\,\alpha_{CD}+\beta_{AB}\,\beta_{CD})\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}\,+\,(\alpha_{AB}\,\beta_{CD}+\beta_{AB}\,\alpha_{CD})\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ (36)
@@ A.2 Generalization to N-Qubit Networks
[…]
@@
Performing a BSM on $(D,E)$, the resulting state of $(A,F)$ has fidelity:
$|\rho_{ADEF}=\big(\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}\big)\,{|{\phi^{+}}\rangle}\bra{\phi^{+}}\,+\,\big(\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}\big)\,{|{\phi^{-}}\rangle}\bra{\phi^{-}}$ (40)
@@ A.3 Imperfect Bell State Measurement
[…]
@@
$\rho$ $|]\displaystyle=\lambda_{BSM}\,\Big[(\alpha_{AB}\,\alpha_{CD}+\beta_{AB}\,\beta_{CD}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AB}\,\beta_{CD}+\beta_{AB}\,\alpha_{CD}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big$] (43)
$|\displaystyle\quad+(1-\lambda_{BSM})\,\Big[\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$
$|]\displaystyle\hskip 85.35826pt+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big$]
@@
[…]
@@
$\rho$ $|]\displaystyle=\lambda_{BSM}^{\prime}\,\Big[(\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big$] (46)
$|\displaystyle\quad+\,(1-\lambda_{BSM}^{\prime})\,\Big[\alpha_{AD}\,\alpha_{EF}{|{++}\rangle}\bra{++}+\alpha_{AD}\,\beta_{EF}{|{+-}\rangle}\bra{+-}$
$|]\displaystyle\hskip 85.35826pt+\beta_{AD}\,\alpha_{EF}{|{-+}\rangle}\bra{-+}+\beta_{AD}\,\beta_{EF}{|{--}\rangle}\bra{--}\Big$]
@@
$\rho_{2}\otimes\rho_{EF}$ $|\displaystyle=\,\Big[\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$ (47)
$|]\displaystyle+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big$]
[…]
@@
$\rho$ $|]\displaystyle=\lambda_{BSM}\,\lambda_{BSM}^{\prime}\,\Big[(\alpha_{AD}\,\alpha_{EF}+\beta_{AD}\,\beta_{EF}){|{\phi^{+}}\rangle}\bra{\phi^{+}}+(\alpha_{AD}\,\beta_{EF}+\beta_{AD}\,\alpha_{EF}){|{\phi^{-}}\rangle}\bra{\phi^{-}}\Big$] (48)
$|\displaystyle\quad+\lambda_{BSM}\,(1-\lambda_{BSM}^{\prime})\,\Big[\alpha_{AD}\,\alpha_{EF}{|{++}\rangle}\bra{++}+\alpha_{AD}\,\beta_{EF}{|{+-}\rangle}\bra{+-}$
$|]\displaystyle\hskip 85.35826pt+\beta_{AD}\,\alpha_{EF}{|{-+}\rangle}\bra{-+}+\beta_{AD}\,\beta_{EF}{|{--}\rangle}\bra{--}\Big$]
$|\displaystyle\quad+(1-\lambda_{BSM})\,\Big[\alpha_{AB}\,\alpha_{CD}{|{++}\rangle}\bra{++}+\alpha_{AB}\,\beta_{CD}{|{+-}\rangle}\bra{+-}$
$|]\displaystyle\hskip 85.35826pt+\beta_{AB}\,\alpha_{CD}{|{-+}\rangle}\bra{-+}+\beta_{AB}\,\beta_{CD}{|{--}\rangle}\bra{--}\Big$]